OpenStudy (anonymous):

Simplify the expression. (3ax/a^2)^4

zepdrix (zepdrix):

Hey book :) Exponents are a little tricky. When it's being applied to a group of stuff, it has to be applied to each things individually.

zepdrix (zepdrix):

Example:\[\large\rm \left(\frac{ab}{c}\right)^5=\frac{a^5b^5}{c^5}\]

OpenStudy (anonymous):

Is this right? \[\left( \frac{ 3^{4}a ^{4}x ^{4} }{ a ^{6} } \right)\]

OpenStudy (anonymous):

@zepdrix

OpenStudy (jdoe0001):

close, just notice that the denominator is \(a^2\) thus \(\bf \left( \cfrac{3ax}{a^2}\right)^4\implies \left( \cfrac{3^1a^1x^1}{a^2}\right)^{\color{brown}{ 4}}\implies \cfrac{3^{1\cdot {\color{brown}{ 4}}}a^{1\cdot {\color{brown}{ 4}}}x^{1\cdot {\color{brown}{ 4}}}}{a^{2\cdot {\color{brown}{ 4}}}} \\ \quad \\ \cfrac{3^4\cancel{a^4} x^4}{\cancel{a^8}} \implies \cfrac{81x^4}{a^2}\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

actually.... hold one sec... darn.. got one mistake there :/

OpenStudy (jdoe0001):

\(\bf \cfrac{3^4\cancel{a^4} x^4}{\cancel{a^8}} \implies \cfrac{81x^4}{a^4}\)

OpenStudy (jdoe0001):

another way to check it, would be \(\bf \cfrac{3^4a^4 x^4}{a^8} \implies 3^4a^4a^{{\color{red}{ -8}}}x^4\implies 3^4a^{4{\color{red}{ -8}}}x^4 \\ \quad \\ 3^4a^{-4}x^4\implies 3^4\cdot \cfrac{1}{a^4}\cdot x^4\implies \cfrac{81x^4}{a^4}\)